OpenStudy (need.answers):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part A: Write functions to represent Cory and Roger's collections throughout the years. Part B: How many cars does Cory have after 6 years? How many does Roger have after the same number of years? Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically.

OpenStudy (sshayer):

for Cory cars after x years\[=15\left( 1+\frac{ 20 }{ 100 } \right)^x=15(1.20)^x\] for Rogers=40+1x=40+2x (after x years) after 6 years for put x=6 C. \[15\left( 1.2 \right)^x=40+x\] by hit and trial method x=6 nearly

OpenStudy (need.answers):

I don't understand the part you put that's dealing with Roger.

OpenStudy (need.answers):

Like 40+1x=40+2x?

OpenStudy (sshayer):

40+1=41 40+2=42 .................. 40+6+46

OpenStudy (sshayer):

correction 40+6=46

OpenStudy (sshayer):

it is wrong typed ,it is 40+x not 40+2x

OpenStudy (need.answers):

So 40+2x is the function to represent Roger?

OpenStudy (need.answers):

40+x *

OpenStudy (sshayer):

no,it is 40+x

OpenStudy (need.answers):

Ah okay. So to continue this and solve the two function, I have to place 6 in x?